OpenStudy (anonymous):

Solve for u. 20 + 17u = -6 + 15u Simplify your answer as much as possible.

OpenStudy (anonymous):

Add 6 on both sides.

OpenStudy (anonymous):

I think i made a mistake.

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Substract 20 on both sides.

OpenStudy (anonymous):

17u=-26+15 l-15 2u=-26 l/2 u=-13

OpenStudy (anonymous):

It is right this time.

OpenStudy (anonymous):

thanks for showing the process